# Koala-36M : A Large-scale Video Dataset Improving Consistency between Fine-grained Conditions and Video Content

### v1

@@ Abstract
As visual generation technologies continue to advance, the scale of video datasets has expanded rapidly, and the quality of these datasets is critical to the performance of video generation models. We argue that temporal splitting, detailed captions, and video quality filtering are three key factors that determine dataset quality. However, existing datasets exhibit various limitations in these areas. To address these challenges, we introduce Koala-36M , a large-scale, high-quality video dataset featuring accurate temporal splitting, detailed captions, and superior video quality. The core of our approach lies in improving the consistency between fine-grained conditions and video content. Specifically, we employ a linear classifier on probability distributions to enhance the accuracy of transition detection, ensuring better temporal consistency. We then provide structured captions for the splitted videos, with an average length of 200 words, to improve text-video alignment. Additionally, we develop a Video Training Suitability Score (VTSS) that integrates multiple sub-metrics, allowing us to filter high-quality videos from the original corpus. Finally, we incorporate several metrics into the training process of the generation model, further refining the fine-grained conditions. Our experiments demonstrate the effectiveness of our data processing pipeline and the quality of the proposed Koala-36M dataset. Our dataset and code will be released at https://koala36m.github.io/.
@@ 1 INTRODUCTION
Generative AI, particularly video generation tasks, has recently garnered significant interest from researchers. These tasks involve generating high-quality videos from textual descriptions or images. A critical factor in the success of these models is the quality of the datasets used for training. Several open-source datasets (e.g. Panda-70M(Chen et al., 2024b), MiraData (Ju et al., 2024), OpenVid (Nan et al., 2024), and VidGen (Tan et al., 2024)) have been introduced, each carefully selecting data sources and applying various evaluation metrics for video filtering. Moreover, innovative approaches have been employed in the video captioning process, such as the multi-modal caption model (Chen et al., 2024b) or structured captions (Ju et al., 2024).
@@
First, the alignment between text and video semantics is essential. Unlike video question answering tasks, where captions are primarily driven by specific question-based details, video generation requires captions that are directly tied to the visual content itself. Due to the infinite granularity of visual signals, this necessitates captions that are rich and detailed. Furthermore, raw video data often contains complex transitions, adding additional challenges in ensuring the accuracy of captions.
@@
Second, the effective evaluation and filtering of low-quality data remains underexplored. Low-quality video data, such as poor visual quality or excessive artificial effects, can impede the training process. However, accurately assessing and filtering such data presents an ongoing challenge. Existing methods typically rely on manually selected quality metrics and heuristic threshold-based filtering, which are often designed for other tasks and may not align with the specific requirements of video generation. As a result, these approaches may not effectively ensure the desired data quality for training.
@@
Third, even with data filtering processes in place, the videos within the dataset still vary in quality, with each video potentially exhibiting different strengths and weaknesses (e.g., one video may have lower clarity but better aesthetic appeal). Training with such heterogeneous data in the same manner may introduce ambiguity for the model, hindering its ability to learn effectively.
@@
To address these issues, we present Koala-36M , a large-scale high-quality video dataset with more accurate video splitting, detailed captions, better data filtering methods and metric conditions. As video content reaches considerable quality, the consistency between fine-grained conditions and video content determines the performance of generation models. We propose a more refined data processing pipeline based on this key insight. Since accurate video splitting leads to better temporal consistency, we first employ a linear classifier on probability distributions to enhance the accuracy of transition detection. Then We generate structured captions for the splitted video clips, with an average length of 200 words, to improve text-video alignment. Sequentially, to prevent the erroneous deletion of high-quality data during filtering, we train a network to predict Video Training Suitability Score (VTSS) on human-aligned datasets to model the joint distribution of sub-metrics. This network takes videos and sub-metrics as input, and outputs a single value called Video Training Suitability Score as the only metric to filter data. Additionally, we introduce data metrics as extra conditions (Metric Conditions) into the generation model during training, helping model distinguish data with different quality and further improving the consistency between fine-grained conditions and video content, which results in better performance and controllability of the generation model.
@@
To further validate Koala-36M and our data processing pipeline, we train video generation models on different datasets. Both the dataset benchmark and the performance of the video generation model demonstrate the advantage of the Koala-36M dataset. We perform more ablation studies to demonstrate effectiveness of our data processing pipeline.
@@ 2 RELATED WORK
Recent advancements in diffusion models have driven the evolution of image generation models into video generation models. In the field of text-to-video (T2V) generation, significant efforts have been made to develop large-scale T2V models, trained on extensive datasets using traditional U-Net-based diffusion architectures (Zeng et al., 2024; Clark & Jaini, 2024; Ge et al., 2023; Yu et al., 2023; Khachatryan et al., 2023) and Transformer-based (DiT) architectures (Ma et al., 2024; Chen et al., 2023b; Lu et al., 2023; Chen et al., 2024a; Xing et al., 2024). The success of these video generation models heavily depends on the quality of the video-text datasets.
@@ 2.1 Video datasets
While several video datasets (Caba Heilbron et al., 2015; Anne Hendricks et al., 2017; Rohrbach et al., 2015; Zhou et al., 2018; Xu et al., 2016; Wang et al., 2023b; Sanabria et al., 2018; Wang et al., 2023a; Chen et al., 2023a) have been applied to tasks such as action recognition, video understanding, visual question answering (VQA), and video retrieval, there remains an urgent need for a high-quality, open-source dataset specifically tailored for training video generation models, providing rich video-text pairs. Datasets such as YouCook2 (Zhou et al., 2018), VATEX (Wang et al., 2019), and ActivityNet (Caba Heilbron et al., 2015) offer high-quality human caption annotations. Another set of datasets, including Miradata (Ju et al., 2024), VidGen-1M (Tan et al., 2024), and OpenVid-1M (Nan et al., 2024), automatically generate high-quality captions and filter data using manually selected thresholds on multiple dataset metrics.
However, these datasets are insufficient in size to support the training of large models. Datasets, including YT-Temporal-180M (Zellers et al., 2021), HD-VILA-100M (Xue et al., 2022), ACAV (Lee et al., 2021), etc., contain hundreds of millions of video-text pairs, but their captions are automatically generated via speech recognition, leading to subpar quality. Panda70M (Chen et al., 2024b), the largest publicly accessible video-text dataset, has become a popular choice for video generation due to its scale and considerable quality. However, its quality still needs further improvement. Specifically, the captions in Panda-70M often provide simplistic, incomplete descriptions of video content, and the frequent transitions in the training videos can result in semantic inconsistencies, potentially leading to undesired or uncertain transitions in the generated videos.
@@ 2.2 Video data curation
As models continue to scale up in size, effective data curation is of paramount importance (Zhou et al., 2023), particularly in the formulation of a well-suited training dataset. This is crucial for enhancing model performance and improving training efficiency during both the pretraining and supervised fine-tuning phases. In the realm of large language models (LLMs), various data curation approaches have been proposed (Xie et al., 2023; Maharana et al., 2023; Tirumala et al., 2023), including optimizations for data quantity, data quality, and domain composition. However, there remains a lack of work exploring data curation strategies in the video domain. Stable Video Diffusion (Blattmann et al., 2023) offers a comprehensive overview of the curation of large-scale video datasets, including techniques such as video clipping, captioning, and filtering. However, the dataset is not open-source. In this study, we propose a novel data processing pipeline for video data and introduce a new video filtering metric. Unlike traditional video quality assessment models (Wu et al., 2023; Zhao et al., 2023; Wu et al., 2022; Sun et al., 2024), which focus primarily on the aesthetic and technical qualities of a video, our approach emphasizes the suitability of videos as training data.
@@ 3 Koala-36M DATASET
Dataset #Videos ATL(words) TVL(hours) Text Filtering Resolution
LSMDC (Rohrbach et al., 2015) 118K 7.0 158 Manual Sub-metrics 1080p
DiDeMo (Anne Hendricks et al., 2017) 27K 8.0 87 Manual Sub-metrics -
YouCook2 (Zhou et al., 2018) 14K 8.8 176 Manual Sub-metrics -
ActivityNet (Caba Heilbron et al., 2015) 100K 13.5 849 Manual Sub-metrics -
MSR-VTT (Xu et al., 2016) 10K 9.3 40 Manual Sub-metrics 240p
VATEX (Wang et al., 2019) 41K 15.2 $\sim$115 Manual Sub-metrics -
WebVid-10M (Bain et al., 2021) 10M 12.0 52K Alt-Text Sub-metrics 360p
HowTo100M (Miech et al., 2019) 136M 4.0 135K ASR Sub-metrics 240p
HD-VILA-100M (Xue et al., 2022) 103M 17.6 760.3K ASR Sub-metrics 720p
VidGen (Tan et al., 2024) 1M 89.3 - Generated Sub-metrics 720p
MiraData (Ju et al., 2024) 330K 318.0 16K Generated & Struct Sub-metrics 720p
Panda-70M (Chen et al., 2024b) 70M 13.2 167K Generated Sub-metrics 720p
Koala-36M (Ours) 36M 202.1 172K Generated & Struct Expert Model 720p
@@ 4 METHOD
As shown in Fig. 3, we propose a refined data processing pipeline for Koala-36M dataset. Our pipeline aims to further improve the consistency between fine-grained conditions and video content. Our main contributions are shown in the red box of Fig. 3. Specifically, we start from the same raw data with Panda-70M (Chen et al., 2024b) dataset. First, we propose a more accurate and efficient transition detection method for video splitting in section 4.1. Then we caption splitted videos with an average length of 200 words based on our structured caption system in section 4.2. Subsequently, we train a Video Training Suitability Score (VTSS) for data filtering to prevent high-quality data from the erroneous deletion in section 4.3. Finally, we introduce multiple data sub-metrics as Metric Conditions into the generation model to enrich the fine-grained conditions in section 4.4.
Figure 3: The proposed data processing pipeline. Compared with previous pipeline, we propose better splitting methods, structured caption system, training suitability assessment network and fine-grained conditioning in red box, improving the consistency between conditions and video content.
@@ 4.1 Video splitting
Splitting videos into temporal splits is crucial for creating video generation datasets. Transition-free video data enable more accurate alignment between text and video, while reducing the difficulty of model training and improving the temporal consistency of generated results. Current video splitting methods Pyscenedetect (Castellano, ) typically detect transitions based on changes in image features between consecutive frames, relying on manually adjusted thresholds as criteria, but often overlook temporal information. As a result, these methods struggle to distinguish between gradual transitions and fast-motion scenes, leading to missed detections in the former and incorrect detections in the latter.
To address the above issues, we first propose a Color-Struct SVM (CSS) module that adopting a learning-based approach for more accurate detection of changes between frames compared to threshold-basd method. Then we leverage temporal smoothing and statistical features to differentiate between gradual transitions and fast-motion scenes.
@@
Then an SVM classifier is employed, using color distance $d_{color}$ and structural distance $d_{struct}$ as the relevant input features; see Eq. 1, Eq. 2, Eq. 3, Eq. 4 . Regarding temporal information, we hypothesize that video changes are relatively stable over time. By estimating a Gaussian distribution of changes from past frames, if the current frame’s change exceeds the $3\sigma$ confidence interval, we consider it a significant transition. This method enhances the differentiation between gradual transitions and fast-motion scenes without increasing computational load. Extensive experiments demonstrate the effectiveness of the transition detection method in  A.1.
@@ 4.2 Video captioning
Detailed captions usually lead to better text-video consistency, which largely determines the granularity of semantic responses. To obtain more detailed captions, we propose a structured caption system, which consists of: (1) the subject, (2) actions of the subject, (3) the environment in which the subject is located, (4) the visual language including style, composition, lighting, etc. (5) the camera language including camera movement, angles, focal length, shot sizes, etc. (6) world knowledge. We generate these aspects separately, and merge them as the final caption.
@@
Similar to previous works (Chen et al., 2024b; Tan et al., 2024; Ju et al., 2024), we first collect a caption dataset by using GPT-4V (OpenAI, 2023) to generate video captions based on our structured system. We then fine-tune a caption model based on LLaVA (Liu et al., 2023) for the entire dataset. Our experiments during fine-tuning show that training the vision encoder improves the accuracy of the caption. And a high-resolution vision encoder helps the caption model capture video details better. To alleviate the computational burden caused by high-resolution inputs, we perform average pooling with a 2x2 kernel on the spatial dimensions of the tokens, ensuring minimal information loss. Notably, we adopt a mixed training strategy involving both static images and dynamic videos, enabling the model to concurrently learn visual understanding in both static and dynamic scenarios. This also enhances data diversity, alleviating the issue of insufficient training samples when solely relying on video data.
When utilizing the caption model to describe videos, a structured caption system often generates longer captions (over 300 words). Different from MiraData (Ju et al., 2024), we limit the caption length to around 200 words. Because the information entropy of the video is finite, and longer captions may repeat mentioned concepts frequently, making it harder for the generation model to extract key information. Finally, we run our captioner on the whole dataset, and the distribution of caption lengths is shown in Fig. 4. Furthermore, we evaluate the quality of captions with caption accuracy and completeness. As shown in Fig. 2 and Tab. 1, our structured caption system significantly improve the quality of captions, providing better text-video consistency.
@@ 4.3 Data filtering
In the large-scale raw dataset, the quality of video content vary significantly. When the performance of the generation model is built upon videos with considerable content quality, it is necessary and crucial to filter out low-quality data and remain high-quality data accurately. Traditional methods often use various sub-metrics to evaluate video quality and then manually set thresholds to filter the desired data. Since these sub-metrics are not completely orthogonal with each other, the video quality is actually a joint distribution of all sub-metrics, which means these thresholds should have implicit constraints with each other. However, existing methods neglect the joint distribution of sub-metrics, resulting in inaccurate thresholds. Meanwhile, since multiple thresholds need to be set, the cumulative effect of inaccurate threshold lead to larger deviations during filtering. Therefore, not only low-quality videos are not correctly filtered out as shown in Fig. 1, but also high-quality videos are mistakenly deleted as shown in Fig. 5.
@@
To address this issue, we propose a Training Suitability Assessment Network (TSA) to model the joint distribution of sub-metrics. This network takes videos and sub-metrics as input, and outputs a single value called Video Training Suitability Score (VTSS) as the only metric to filter data. This score reflects whether a video is suitable for training purposes. Specifically, we first collect the training set from human evaluation based on a new criteria. Then we train the Training Suitability Assessment Network (TSA) and employ it to calculate VTSS for all videos. Finally, we set a single threshold for VTSS based on its distribution to filter desired data.

### v2

@@ Abstract
With the continuous progress of visual generation technologies, the scale of video datasets has grown exponentially. The quality of these datasets plays a pivotal role in the performance of video generation models. We assert that temporal splitting, detailed captions, and video quality filtering are three crucial determinants of dataset quality. However, existing datasets exhibit various limitations in these areas. To address these challenges, we introduce Koala-36M , a large-scale, high-quality video dataset featuring accurate temporal splitting, detailed captions, and superior video quality. The essence of our approach lies in improving the consistency between fine-grained conditions and video content. Specifically, we employ a linear classifier on probability distributions to enhance the accuracy of transition detection, ensuring better temporal consistency. We then provide structured captions for the splitted videos, with an average length of 200 words, to improve text-video alignment. Additionally, we develop a Video Training Suitability Score (VTSS) that integrates multiple sub-metrics, allowing us to filter high-quality videos from the original corpus. Finally, we incorporate several metrics into the training process of the generation model, further refining the fine-grained conditions. Our experiments demonstrate the effectiveness of our data processing pipeline and the quality of the proposed Koala-36M dataset. Our dataset and code have been released at https://koala36m.github.io/.
@@ 1 Introduction
Generative AI, especially in video generation tasks, has recently captivated the attention of researchers. These tasks entail generating high - quality videos from textual descriptions or images. The quality of the datasets used for training is a decisive factor in the success of these models. Multiple open - source datasets, such as Panda - 70M[10], MiraData [14], OpenVid [22], and VidGen [27], have been introduced. Each of these datasets carefully selects data sources and applies diverse evaluation metrics for video filtering. Additionally, innovative methods, like the multi - modal caption model [10] or structured captions [14], have been utilized in the video captioning process.
@@
First, the alignment between text and video semantics is essential. Unlike video question - answering tasks, where captions are primarily driven by specific question - based details, video generation requires captions that are directly tied to the visual content itself. Given the infinite granularity of visual signals, this demands captions that are rich and detailed. Moreover, raw video data often contains complex transitions, posing additional challenges in ensuring caption accuracy. Second, the effective evaluation and filtering of low - quality data remain underexplored. Low - quality video data, such as those with poor visual quality or excessive artificial effects, can impede the training process. However, accurately assessing and filtering such data presents an ongoing challenge. Existing methods typically rely on manually selected quality metrics and heuristic threshold - based filtering, which are often designed for other tasks and may not meet the specific requirements of video generation. Consequently, these approaches may not effectively guarantee the desired data quality for training.
@@
Third, even with data filtering processes in place, the videos within the dataset still vary in quality, with each video potentially having different strengths and weaknesses (e.g., one video may have lower clarity but better aesthetic appeal). Training with such heterogeneous data in the same way may introduce ambiguity for the model, hampering its ability to learn effectively.
@@
To address these issues, we present Koala-36M , a large-scale high-quality video dataset with more accurate video splitting, detailed captions, better data filtering methods and metric conditions. As video content reaches considerable quality, the consistency between fine-grained conditions and video content determines the performance of generation models. Based on this crucial understanding, we propose a more sophisticated data processing pipeline. Since accurate video splitting leads to better temporal consistency, we first employ a linear classifier on probability distributions to enhance the accuracy of transition detection. Then we generate structured captions for the segmented video clips, with an average length of 200 words, to improve text-video alignment. Moreover, to prevent the accidental elimination of high-quality data during the filtering procedure, we train a network on human - curated datasets to predict the Video Training Suitability Score (VTSS). This network models the joint distribution of sub-metrics by taking videos and sub - metrics as inputs and outputs a single Video Training Suitability Score, which serves as the sole metric for data filtering. Additionally, we introduce data metrics as supplementary conditions (Metric Conditions) into the generation model during training. This enables the model to differentiate data of varying quality levels and further enhances the consistency between fine - grained conditions and video content, which results in better performance and controllability of the generation model.
@@
To further validate the efficacy of Koala-36M and our data processing pipeline, we conduct training of video generation models on different datasets. Both the dataset benchmark and the performance of the video generation model demonstrate the advantage of the Koala-36M dataset. We perform more ablation studies to demonstrate effectiveness of our data processing pipeline.
@@ 2 Related Work
Recent progress in diffusion models has spurred the evolution of image generation models into video generation models. In the realm of text-to-video (T2V) generation, considerable efforts have been exerted to develop large-scale T2V models. These models are trained on extensive datasets, making use of traditional U-Net-based diffusion architectures [41, 11, 12, 38, 15] and Transformer - based (DiT) architectures [19, 8, 18, 9, 35]. The success of these video generation models is critically reliant on the quality of the video-text datasets.
@@ 2.1 Video datasets
While several video datasets [5, 1, 24, 44, 36, 31, 25, 29, 7] have been applied to tasks such as action recognition, video understanding, visual question answering (VQA), and video retrieval, there remains an urgent need for a high-quality, open-source dataset specifically tailored for training video generation models, providing rich video-text pairs. Datasets such as YouCook2 [44], VATEX [30], and ActivityNet [5] offer high-quality human caption annotations. Another set of datasets, including Miradata [14], VidGen-1M [27], and OpenVid-1M [22], automatically generate high-quality captions and filter data using manually selected thresholds on multiple dataset metrics.
However, these datasets are insufficient in size to support the training of large models. Datasets, including YT-Temporal-180M [40], HD-VILA-100M [37], ACAV [16], etc., contain hundreds of millions of video-text pairs, but their captions are automatically generated via speech recognition, leading to subpar quality. Panda70M [10], the largest publicly available video - text dataset, has gained popularity for video generation tasks owing to its scale and relatively good quality. Nevertheless, its quality still demands further enhancement. In particular, the captions in Panda - 70M frequently offer simplistic and incomplete descriptions of video content. Moreover, the frequent transitions in the training videos can cause semantic incoherence, potentially giving rise to undesirable or unpredictable transitions in the generated videos.
Dataset #Videos ATL(words) TVL(hours) Text Filtering Resolution
LSMDC [24] 118K 7.0 158 Manual Sub-metrics 1080p
DiDeMo [1] 27K 8.0 87 Manual Sub-metrics -
YouCook2 [44] 14K 8.8 176 Manual Sub-metrics -
ActivityNet [5] 100K 13.5 849 Manual Sub-metrics -
MSR-VTT [36] 10K 9.3 40 Manual Sub-metrics 240p
VATEX [30] 41K 15.2 $\sim$115 Manual Sub-metrics -
WebVid-10M [2] 10M 12.0 52K Alt-Text Sub-metrics 360p
HowTo100M [21] 136M 4.0 135K ASR Sub-metrics 240p
HD-VILA-100M [37] 103M 17.6 760.3K ASR Sub-metrics 720p
VidGen [27] 1M 89.3 - Generated Sub-metrics 720p
MiraData [14] 330K 318.0 16K Generated & Struct Sub-metrics 720p
Panda-70M [10] 70M 13.2 167K Generated Sub-metrics 720p
Koala-36M (Ours) 36M 202.1 172K Generated & Struct Expert Model 720p
@@ 2.2 Video data curation
As models continuously expand in scale, effective data curation is of paramount importance [43], particularly in the formulation of a well-suited training dataset. This is crucial for enhancing model performance and improving training efficiency during both the pretraining and supervised fine-tuning phases. In the realm of large language models (LLMs), various data curation approaches have been proposed [34, 20, 28], including optimizations for data quantity, data quality, and domain composition. However, research on exploring data curation strategies within the video domain remains scarce. Stable Video Diffusion [3] offers a comprehensive review of the curation of large - scale video datasets, encompassing techniques such as video clipping, captioning, and filtering. Regrettably, the dataset is not open-source. In this study, we propose a novel data processing pipeline for video data and introduce a new video filtering metric. Distinct from traditional video quality assessment models [33, 42, 32, 26], which mainly concentrate on the aesthetic and technical aspects of a video, our approach places emphasis on the suitability of videos for training purposes.
@@ 4 Method
As shown in Fig. 3, we propose a refined data processing pipeline for Koala-36M dataset. Our pipeline aims to further improve the consistency between fine-grained conditions and video content. Our main contributions are highlighted in the red box of Fig. 3. Specifically, we start from the same raw data with Panda-70M [10] dataset. Firstly, in section 4.1, we propose a more precise and efficient transition detection approach for video segmentation. Subsequently, in section 4.2, we generate captions for the segmented videos, with an average length of 200 words, using our structured caption system. Next, in section 4.3, we train a Video Training Suitability Score (VTSS) for data filtering, aiming to prevent the inadvertent deletion of high - quality data. Finally, in section 4.4, we incorporate multiple data sub - metrics as Metric Conditions into the generation model to enrich the fine - grained conditions.
Figure 3: The proposed data processing pipeline. Compared with previous pipeline, we propose better splitting methods, structured caption system, training suitability assessment network and metrics conditioning in red box, improving the consistency between conditions and video content.
@@ 4.1 Video splitting
Splitting videos into temporal segments is crucial for creating video generation datasets. Transition-free video data enable more accurate alignment between text and video, while reducing the difficulty of model training and improving the temporal consistency of generated results. Existing video splitting techniques Pyscenedetect [6] typically detect transitions based on changes in image features between consecutive frames, relying on manually adjusted thresholds as criteria, but often overlook temporal information. As a result, these methods struggle to distinguish between gradual transitions and fast-motion scenes, leading to missed detections in the former and incorrect detections in the latter.
To tackle the aforementioned problems, we initially introduce a Color-Struct SVM (CSS) module. This module employs a learning-based strategy, enabling more precise detection of frame-to-frame changes compared to threshold-based methods. Subsequently, we utilize temporal smoothing and statistical features to distinguish between gradual transitions and fast-motion scenes.
@@
Then an SVM classifier is employed, using color distance $d_{color}$ and structural distance $d_{struct}$ as the relevant input features; see Eq. 1, Eq. 2, Eq. 3, Eq. 4 . Regarding temporal information, we hypothesize that video changes are relatively stable over time. By estimating a Gaussian distribution of changes from past frames, if the current frame’s change exceeds the $3\sigma$ confidence interval, we consider it a significant transition. This method enhances the differentiation between gradual transitions and fast-motion scenes without increasing computational load. Extensive experiments demonstrate the effectiveness of the transition detection method in  A.
@@ 4.2 Video captioning
Detailed captions typically result in enhanced text - video consistency, which significantly influences the granularity of semantic responses. To obtain more detailed captions, we propose a structured caption system, which consists of: (1) the subject, (2) actions of the subject, (3) the environment in which the subject is located, (4) the visual language including style, composition, lighting, etc. (5) the camera language including camera movement, angles, focal length, shot sizes, etc. (6) world knowledge. We generate these aspects separately, and merge them as the final caption.
@@
Similar to previous works [10, 27, 14], we first collect a caption dataset by using GPT-4V [23] to generate video captions based on our structured system. We then fine-tune a caption model based on LLaVA [17] for the entire dataset. Our experiments during fine-tuning show that training the vision encoder improves the accuracy of the caption. And a high-resolution vision encoder helps the caption model capture video details better. To alleviate the computational burden caused by high-resolution inputs, we perform average pooling with a 2x2 kernel on the spatial dimensions of the tokens, ensuring minimal information loss. Notably, we implement a mixed training strategy involving both static images and dynamic videos, enabling the model to concurrently learn visual understanding in both static and dynamic scenarios. This also enhances data diversity, alleviating the issue of insufficient training samples when solely relying on video data.
Finally, we run our captioner on the whole dataset, and the distribution of caption lengths is shown in Fig. 4. Furthermore, we evaluate the quality of captions with caption accuracy and completeness. As shown in Fig. 2 and Tab. 1, our structured caption system significantly improve the quality of captions with better text-video consistency.
@@ 4.3 Data filtering
In the large-scale raw dataset, the quality of video content varies significantly. When the performance of the generation model hinges on videos with considerable content quality, it is both necessary and crucial to filter out low-quality data and retain high-quality data accurately. Traditional approaches typically utilize various sub-metrics to evaluate video quality and then manually set thresholds to filter the desired data. Since these sub-metrics are not completely orthogonal with each other, the video quality is actually a joint distribution of all sub-metrics, which means these thresholds should have implicit constraints with each other. However, existing methods neglect the joint distribution of sub-metrics, resulting in inaccurate thresholds. Meanwhile, since multiple thresholds need to be set, the cumulative effect of inaccurate threshold leads to larger deviations during filtering. Therefore, not only low-quality videos are not correctly filtered out in Fig. 1, but also high-quality videos are mistakenly deleted in Fig. 5. More analysis and experiments are shown in App. D.
@@
To address this issue, we propose a Training Suitability Assessment Network (TSA) to model the joint distribution of sub-metrics. This network takes videos and sub-metrics as input, and outputs a single value called Video Training Suitability Score (VTSS) as the only metric to filter data. This score reflects whether a video is suitable for training purposes. Specifically, we first collect the training set from human evaluation based on a new criteria. Then we train the Training Suitability Assessment Network (TSA) and employ it to calculate VTSS for all videos. Finally, we set a single threshold for VTSS based on its distribution for filtering.